OpenStudy (anonymous):

forgot this one: solve for the unknown. 6=7ke^0.5

OpenStudy (anonymous):

my working:\[\frac{ 6 }{ e ^{0.5}} =7k\]

OpenStudy (anonymous):

is that fine so far?

OpenStudy (anonymous):

yes , i think u got a solution for k there

OpenStudy (anonymous):

@waterineyes :)

OpenStudy (anonymous):

okay then i simplify 6/e^0.5 and then divide by 7?

OpenStudy (anonymous):

yes :)

OpenStudy (saifoo.khan):

\[\frac{6}{e^{0.5} \times 7} = k\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

welcome